OpenStudy (anonymous):

What is the sum of a 10–term arithmetic sequence where the first term is 3 and the last term is 75?

OpenStudy (mertsj):

\[\S _{10}=\frac{1}{2}(10)(3+75))=5(78)=390\]

OpenStudy (anonymous):

first term is 3 ,so a=3 and last term is 75 ,so L=75 _________ sn=n\2(a+L) ,s10=10\2(3+75)=390